OpenStudy (anonymous):

I need help converting these polar coordinates into rectangular coordinates. (2,7pi/3)

terenzreignz (terenzreignz):

Luckily, polar to rectangular is far easier than the other way around :D If you have the polar coordinates \[\large (\color{green}r,\color{orange}\theta)\] then their corresponding pair of rectangular coordinates is given by \[\Huge (\color{red}x,\color{blue}y) = (\color{green}r\cos(\color{orange}\theta) \ , \ \color{green}r\sin(\color{orange}\theta))\]

OpenStudy (anonymous):

x=rcosA y=rsinA square and addboth equations this gives x^2 + y^2 = r^2 r = sqrt(x^2 + y^2) now have r and x and put them in equations and then solve. if any doubt , can ask furthur :)

OpenStudy (anonymous):

Thnk you everyone

OpenStudy (anonymous):

doesn't matter